OpenStudy (anonymous):

If the volume of a sphere is 10293.33 cm3, what is the diameter? Use 3.14 to approximate π. Round answers to the nearest hundredth. Question 18 options: a) 13.50 cm b) 27 cm c) 31.4 cm d) 13.9 cm

OpenStudy (skullpatrol):

You will need to know the formula for the volume of a sphere.

OpenStudy (anonymous):

\[V=\frac{ 4 }{ 3 }\pi r ^{3}\] I'm back early

OpenStudy (anonymous):

ok so what is the formula

OpenStudy (anonymous):

o I dint see that ok I have the formula what do I do now

OpenStudy (skullpatrol):

How are r and d related?

OpenStudy (skullpatrol):

r=radius and d=diameter

OpenStudy (anonymous):

well we're going to reverse it \[10293.33=\frac{ 4 }{ 3 }\left( 3.14 \right)r ^{3}\]

OpenStudy (anonymous):

idk how they are related

OpenStudy (anonymous):

ok siegel

OpenStudy (anonymous):

so first you want to get r alone, so what can you do

OpenStudy (anonymous):

10293.33 cm ^3

OpenStudy (anonymous):

10293.33 * ^3

OpenStudy (anonymous):

well here for now let's just look at (4/3)pi, what can you do

OpenStudy (anonymous):

I got 200.96

OpenStudy (anonymous):

i did 4*4*4=64*3.14=200.96

OpenStudy (anonymous):

ok slow down. first let's divide 10293.33 by 3.14. do you know why?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3278.1305732484 was what i got

OpenStudy (anonymous):

good we want to get r alone, so to do this we have to get rid of everything else. first we divide both sides by 3.14 to cancel it out. now what's the next thing to divide.

OpenStudy (anonymous):

3278.1305732484

OpenStudy (anonymous):

right so what does the equation look like now that we've canceled 3.14

OpenStudy (anonymous):

3278.1305732484 /3.14

OpenStudy (anonymous):

well the whole equation so\[3278.1305732484=\frac{ 4 }{ 3 }r^3\]

OpenStudy (anonymous):

do you understand how we got there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well we had \[10293.33=\frac{ 4 }{ 3 }(3.14)r^3\]then we divided 3.14 from both sides\[\frac{ 10293.33 }{ 3.14 }=\frac{ \frac{ 4 }{ 3 }(3.14)r^3 }{ 3.14 }\]the 3.14s cancel on the right side and 10293.33 is divided by 3.14 on the right side\[3278.1305732484=\frac{ 4 }{ 3 }r^3\] does that clear it up?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so how do I get the answer

OpenStudy (anonymous):

well we have \[3278.1305732484=\frac{ 4 }{ 3 }r^3\]what can you divide next to get r alone

OpenStudy (anonymous):

3278.1305732484/3.14

OpenStudy (anonymous):

ok I have about 15 minuets left to finisj

OpenStudy (anonymous):

we can do this, now the 3.14 is gone, but what about that 4/3, what do you think you do with that?

OpenStudy (anonymous):

4/3 divided by 3.14

OpenStudy (anonymous):

close you are right about the 4/3 but we want to cancle it from the right side. so what can we do to the left side with the 4/3

OpenStudy (anonymous):

* it

OpenStudy (anonymous):

8 minutes left

OpenStudy (anonymous):

dont worry, you multiply the reciprocal, do you know what that is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the reciprocal of 4/3 is 3/4, that's how you divide with fractions, you multiply the reciprocal.

OpenStudy (anonymous):

I got 5184

OpenStudy (anonymous):

4*4*4=64*3*3*3*3=5184

OpenStudy (anonymous):

ok I have like 3 minuets to finish this question

OpenStudy (anonymous):

there is no 4, ok look multiply 3278.1305732484 by 3/4, then find the third root. you need a good calculator for this. it will look like\[3278.1305732484\times \frac{ 3 }{ 4 }=r^3\]then like this\[2458.59793=r^3\]then\[\sqrt[3]{2458.59793}=\sqrt[3]{r ^{3}}\]and then finally like this\[r=13.49674897\]if you still need help maybe we can talk it through tomorrow.

OpenStudy (anonymous):

ok so is a or c the answer

OpenStudy (anonymous):

a. do you understand kind of?

OpenStudy (anonymous):

yeah kind of there just 1 last question I don't know the diffre

OpenStudy (anonymous):

Kelly likes to play a game with her friends where she stands in one spot with one end of a jump rope in her hand. She then twirls the rope around on the ground in a circle. Her friends jump over the other end of the rope as it comes near them. In this game, does the rope more closely resemble a radius or a diameter of a circle? radius diameter

OpenStudy (anonymous):

i think it is diameter am i right or wrong

OpenStudy (anonymous):

it's radius